OpenStudy (anonymous):

(10x^6+20x^4-15x^2)/5x^2=_x^4+_x^_ _

OpenStudy (anonymous):

the _'s are blank spaces that i need to fill in

OpenStudy (anonymous):

Slowly divide everything by 5x^2. Start with 10x^6. What does that give you?

OpenStudy (anonymous):

2x^3+4x^2-3x?

OpenStudy (anonymous):

nope. 10x^6 = 10 * x * x * x * x * x *x Divide that by 5x^2 = 5 * x * x You can cancel out the 'x's, so that gives you (10/5) * x * x * x * x, which is 2x^4. Try for the other two.

OpenStudy (anonymous):

2x^4 + 4x^2 - 3 ?

OpenStudy (anonymous):

You got it! :)

OpenStudy (anonymous):

thank you!